Applebee:

A perpendicular bisector, , is drawn through point C on . If the coordinates of point A are (-3, 2) and the coordinates of point B are (7, 6), the x-intercept of is . Point lies on .

surjithayer:

let (x,y) be the mid-point. \[x=\frac{ -3+7 }{ 2 }=\frac{ 4 }{ 2}=2\] \[y=\frac{ 2+6 }{ 2 }=\frac{ 8 }{ 2}=4\] midpoint=(2,4) slope of AB\[=\frac{ 6-2 }{ 7+3 }=\frac{ 4 }{ 10 }=\frac{ 2 }{5 }\] slope of line perpendicular to AB=-5/2 eq. of line pependicular to AB through (2,4) is \[y-4=-\frac{ 5 }{ 2 }(x-2)\] 2y-8=-5x+10 5x+2y=10+8 or 5x+2y=18 perpendicular bisector of AB lies on line 5x+2y=18

Applebee:

thank you